. Acme Corporation has $1,200,000 in assets and $700,000 of debt. It reports net income of $200,000.

  1. A )What is the return on assets?
  2. B) What is the return on stockholder’s equity?
  3. C) If the firm has an asset turnover ratio of 2.5 times, what is the profit margin (return on sales)?
Answer #1
A) Return on assets = Net income / Total assets = 200000 / 1200000 16.67%
B)
Stockholders' equity = Total assets - Total debt = 1200000 - 700000 500000
Return on stockholders' equity = Net income / Stockholders' equity = 200000 / 500000 40%
C)
Asset turnover ratio = Sales / Total assets
2.5 = Sales / 1200000
Sales = 2.5 * 1200000 3000000
Profit margin = Net income / Sales = 200000 / 3000000 6.67%